Use the information in the table below to answer questions 13 through 15. In these exercises, the deaths are included in the discharges; this includes deaths occurring in less than 48 hours and postoperative deaths.

Community Hospital

Annual Statistics 20XX

Service

Discharges

Deaths

< 48 hours

Postop

Medicine

1,478

92

8

0

Surgery

1,385

25

10

2

Obstetrics

751

1

0

0

Newborn

753

1

1

0

Psychiatric

486

2

0

0

Rehabilitation

362

22

1

0

Total

5,215

143

20

2

Additional Information:

Anesthetics administered

1,386

Anesthetic deaths

1

Patients operated on

1,385

Total livebirths

753

Fetal deaths:

Early

5

Intermediate

5

Late

1

13. What is the gross death rate for the Rehabilitation Service?

a. 0.67%

b. 0.68%

c. 6.07%

d. 6.08%

14. What is the total hospital net death rate?

a. 0.24%

b. 2.36%

c. 2.37%

d. 23.68%

15. What is the postoperative death rate?

a. 0.14%

b. 1.44%

c. 14.10%

d. 14.44%
Answer #1

Solution:

13.

Step 1:

The gross death rate of rehabilitative services is the ratio of total deaths of rehabilitative services to the total discharges of rehabilitative services.

Step 2:

Gross death rate of rehabilitative services = Total no. of rehabilitative death / No. of rehabilitative discharges X 100

Gross death rate of rehabilitative services = 22 / 362 X 100

= 6.07%

Step 3:

Hence the option c. 6.07%.

14.

Step 1:

Total Hospital net death rate = (Total no.of death - death under 48 hours) / (total no.of discharges - death under 48 hours) X 100

= (143-20) / (5215 - 20) X 100

= 123 / 5195 X 100

= 0.0236 X 100

= 2.36%

Step 2:

Hence the option b. 2.36% is correct.

15.

Step 1:

Postoperative death rate = Total no.of deaths within 10 days after surgery / Total no.of patients operated in the period X 100

= 2 / 1385 X 100

= 0.144%

Step 2:

Hence the option a.0.14% is correct.